OpenStudy (anonymous):

Three angles A,B and C. Angle A=3times B. Angle B=3times C. Angles A+C=90degrees. What is angle B

OpenStudy (anonymous):

Try to put together an equation with only one unknown. If A = 3B, B = 3C, and A+C = 90, Then A = 3B and 3(3C). So, instead of writing A + C = 90, we can write 3(3C) + C = 90. Now solve that for C and you should have no problem finding B since we know that it is 3 times C

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

yes and thank you very much